OpenStudy (anonymous):

What is the solution to the equation? 2q + 18 = –5q – 3 A. q = –5 B. q = –3 C. q = 3 D. q = 5

OpenStudy (gorv):

2q+5q=-3-18 7q=-21 q=-3

OpenStudy (anonymous):

Thanks for the help! :D